# and quantum information science

### v1

@@ 4.2 Complexity saturation as a late-time stage of quantum many-body equilibration
The second law of thermodynamics suggests that state complexity should grow under generic dynamics. Of course, the second law concerns entropy, not complexity, and originally described classical systems. Yet complexity seems like it ought to obey the second law’s spirit [44]. It does, serving as the final stage of quantum many-body equilibration.
@@
To review the second law, we recall a classical monatomic gas in a box. Suppose that the gas particles are bunched together in a corner. The particles will spread across the box: the gas’s entropy, calculable with the Sackur-Tetrode equation, will end up greater than it was originally. The gas will equilibrate internally. Furthermore, suppose that the particles interact. They will thermalize; a particle’s phase-space distribution will approach the Maxwell–Boltzmann distribution.
@@
Given the second law’s ubiquity, one should expect generic quantum many-body systems, too, to equilibrate and thermalize internally. Quantum many-body systems undergo more stages of equilibration and thermalization than classical systems do. We will synopsize some stages, which end with complexity saturation. First, we discuss the kinds of systems expected to thermalize.
@@
Which systems thermalize is difficult to predict and prove. Thermalization is often attributed to chaos, but many definitions of quantum chaos exist [7, 8, 9]. According to early thinking, a system is quantum-chaotic if it has a chaotic semiclassical limit. Later studies defined chaos through energy-level statistics modeled with Wigner–Dyson distributions. Recent trends define chaos through out-of-time-ordered correlators [45] and spectral form factors [46, 47, 48, 7], reviewed below. Some authors acknowledge the ambiguity by writing “chaos” with quotation marks and meaning nonintegrability [49]. Yet debate surrounds the definition of quantum nonintegrability, too. According to a common definition, a system is integrable if it has extensively many nontrivial conserved quantities [8].888 To understand the nontrivial, consider eigendecomposing an arbitrary quantum Hamiltonian: $H=\sum_{j}E_{j}\lvert j\rangle\!\langle j\rvert$. The Hamiltonian conserves every eigenprojector $\lvert j\rangle\!\langle j\rvert$. Furthermore, the number of eigenprojectors is exponentially large in the system size. Yet these facts do not render $H$ integrable; if they did, every $H$ would be integrable. A conserved quantity must be unlike the $\lvert j\rangle\!\langle j\rvert$s to be nontrivial. To compound the ambiguity, many equilibration results follow from empirical observations, rather than from proofs, and govern specific systems, rather than general ones. We therefore write “chaotic,” with quotation marks, to communicate expectations about behaviors common to many systems called chaotic or nonintegrable.
@@
In reviewing stages of quantum many-body equilibration, we refer to the following setup. Consider a closed quantum system of $N\gg 1$ subsystems—for example, qubits. Let the Hamiltonian $H$ be “chaotic”; and let the interactions be $k$-body, for some $k\in\{1,2,\ldots,N\}$. Suppose that the system begins in a simple, pure nonequilibrium state, such as an $N$-qubit product state.
@@
We now define two quantities—the out-of-time-ordered correlator and spectral form factor—used to define stages of equilibration. Let $W$ and $V$ denote local, Hermitian and/or unitary operators localized far apart. Examples include Pauli operators that act on distant qubits. Define the Heisenberg-picture operator $W(t)\coloneqq e^{iHt}We^{-iHt}$, and consider an arbitrary state $\rho$. The (four-point) out-of-time-ordered correlator is ${\rm Tr}\bm{(}W^{\dagger}(t)V^{\dagger}W(t)V\rho\bm{)}$; we discuss its significance below. Define the partition function $Z(\beta)\coloneqq{\rm Tr}(e^{-\beta H})$. If $\beta,t\in\mathbb{R}$, then the spectral form factor is $|Z(\beta+it)/Z(\beta)|^{2}$. It measures correlations across a Hamiltonian’s spectrum.
@@
    Local equilibration: Two-point correlators decay to $\sim 1/e$ of their initial values. Time-ordered correlators relax to approximately their long-time values. Small subsystems thermalize. This equilibration occurs at the dissipation time $t_{\rm D}$ [5]. $t_{\rm D}$ is also called the thermalization time. It is called the collision time if a system admits of a quasiparticle description. Under certain conditions (e.g., if the interactions are highly nonlocal), $t_{\rm D}$ is constant in $N$. Under other conditions, $t_{\rm D}\sim N$. This latter scaling characterizes a one-dimensional system with nearest-neighbor interactions, by the Lieb–Robinson bound [50].
@@
    Quantum information scrambling: Initially localized information spreads across the system through many-body entanglement. Out-of-time-ordered correlators decay from their $O(1)$ initial values [45]. This scrambling occurs around the scrambling time, $t_{*}$. $t_{*}$ can be parametrically larger than $t_{\rm D}$, as under all-to-all coupling [41]. A lower bound limits how quickly many-body entanglement can form: $t_{*}\geq\frac{\hbar}{2\pi k_{\rm B}T}\,\log(N)$ [5]. This bound saturates only if $t_{\rm D}$ is constant. In the context of semiclassical, single-particle quantum chaos, an analogue of $t_{*}$ is called the Ehrenfest time [51].
@@
    Breakdown of quantum–classical correspondence and rise in predictiveness of random-matrix theory: This stage can begin around the same time as scrambling. Hence we have already seen one name for the time when this stage takes place: the Ehrenfest time $t_{\rm E}$, also called the Thouless time $t_{\rm Th}$. This time scale varies logarithmically with the system size: $t_{\rm E}\sim\log(N)$.
    How this stage manifests depends on the system and measures studied. An example setting involves transport in, e.g., electronic systems [52, 53]. Here, an initially narrow wave packet expands to cover a classically relevant length scale around $t_{\rm E}$. The quantum system’s behavior diverges from its classical analogue’s; hence the name Ehrenfest time alludes to Ehrenfest’s theorem, which highlights a parallel between quantum and classical systems. In another example, at this stage, the spectral form factor changes from dipping to ramping upward: random-matrix theory begins to accurately predict the spectral form factor, which ceases to depend on microscopic details [54, 55].
@@
    The spectral form factor changes from increasing linearly with time to remaining constant: This change happens at the Heisenberg time $t_{\rm H}$, or plateau time [54]. Denote by $\Delta$ the average inverse gap between consecutive eigenenergies. The Heisenberg time is $t_{\rm H}\coloneqq h/\Delta\sim e^{N}$ [53].
@@
    Complexity saturation: The state’s complexity grows to $\sim e^{N}$. The saturation time has been conjectured to scale as $e^{N}$. This conjecture has been proven under certain conditions (Sec. 4.3).
As a caveat, different authors use the same terms differently, when referring to stages of quantum many-body equilibration. Also, the above list is intended to be illustrative, not comprehensive; compiling all the stages [9, 56, 57, 55, 3] would require another review. For example, subtleties in the spectral form factor define other stages; see Sec. 1.1 of [56], and Sec. 1.4.3 of [55], for clear synopses and for references. Operator dynamics in Krylov space require their own treatment [3]. Additionally, if $N$ is finite, then revivals (Poincaré recurrences and complexity revivals) occur long after times $\sim e^{N}$. Such revivals temporarily undo the equilibration but are consistent with the second law, given the system’s finite size.
@@ 4.3 Growth of complexity under random circuits
Random circuits form a toy model for chaotic quantum dynamics. We can prove random-circuit versions of conjectures originally formulated about chaotic dynamics’ complexities. When forming a random circuit, we choose the gates independently from a probability distribution on the unitary group $\mathrm{SU}(4)$. Typically, we choose gates according to the Haar measure $\mu_{\rm H}$, the unique distribution that is invariant under left- and right- multiplication by any unitary. $\mu_{\rm H}$ generalizes the uniform distribution on an interval, as well as the Lebesgue measure, to (locally compact) Lie groups. $\mu_{\rm H}$ can be viewed as the uniform measure on these groups.
Known barriers such as the natural proof barrier discovered by Razborov and Rudich [58, 59, 60] inhibit us from proving superpolynomial lower bounds on the circuit complexity of an explicit Boolean function. The minimal number of elementary gates (such as AND, NAND, and XOR) in a Boolean function’s implementation is called its circuit complexity. Similar obstructions might inhibit us from proving superpolynomial lower bounds on quantum states’ quantum complexities. That is, the minimal number of $2$-local quantum gates in any implementation of a state. One can more easily prove lower bounds on the complexities of states prepared with random circuits as random circuits quickly assume properties of the Haar measure, counting arguments about which yield lower bounds on the circuit complexity. In 1949 Claude Shannon [61] observed that the number of Boolean functions $f:\{0,1\}^{n}\to\{0,1\}$ outgrows the number of functions implementable with a subexponential number of elementary gates. A similar counting argument applies to Haar-random unitaries [62].
Sections 4.4 and Section 4.5 review two results about the growth of quantum circuit complexity. Both arguments followed from adapting Haar-measure counting arguments to random circuits. In a sense, these arguments offer to replace statements about vaguely defined “generic” dynamics as in section 4.1 with well-defined statements about average dynamics (which would intuitively be controlled by the dynamics of most systems and so would be “generic” in a well defined sense). This is done by thinking about which exact statistics/distribution should we consider for the unitaries and the circuits where now generic takes the sense of "very likely" or "of likelihood=1". Section 4.6 reviews why we still lack satisfying lower bounds on the quantum complexity of chaotic dynamics.
@@ 4.4 Linear growth of exact circuit complexity
Here, we review a proof of the exact circuit complexity’s linear growth [63]. More precisely, we show that the exact circuit complexity grows linearly with the number of layers, up to exponentially deep random quantum circuits, with unit probability over the choice of circuits. Moreover, the complexity saturates at exponentially deep circuits, again with unit probability. Just as random circuits form a toy model for chaotic dynamics, the exact circuit complexity forms a toy model for quantum complexity: the exact circuit complexity quantifies the minimal number of gates required to prepare a state or effect a unitary exactly. The preparation and effecting cannot include errors. Consequently, some unitaries arbitrarily close to the identity operator have exact circuit complexities of $e^{\Omega(K)}$. [^9] Throughout the rest of this section, we use big-Omega notation: $\Omega(X)$ means, roughly, “a term that grows at least as quickly as $X$”. Recall that $K$ denotes the system size, or number of degrees of freedom (qubits).
We now sketch a proof of Brown and Susskind’s conjecture [44] that circuit complexity grows linearly for an exponentially long time, as applied to random circuits and the exact circuit complexity. We will compare the degrees of freedom in two sets of unitaries: the unitaries generated by the depth-$D$ circuits and the unitaries generated by smaller, depth-$D^{\prime}$ circuits. The basic idea echoes the fact that, in $\mathbb{R}^{3}$, a plane has no volume. In terms of a well-behaved probability measure on a manifold $M$, every lower-dimensional submanifold has a probability of 00. Denote by $\mathcal{U}_{R}$ the set of all unitaries that can be generated by $R$ gates each. Fortunately, $\mathcal{U}_{R}$, although not a manifold, turns out to be a semialgebraic set defined by constraints $\{f_{i}\geq 0,g_{i}>0,h_{i}=0\}$, where $f_{i},g_{i},h_{i}:{\rm SU}(D)\to\mathbb{R}$ are polynomials. Every semialgebraic set has a well-defined dimension similar to a manifold’s dimensionality.
@@ 4.5 Linear growth of approximate circuit complexity
We can combine this bound with random circuits’ quick convergence to unitary designs. Brown and Viola provided evidence for fast convergence by analyzing the spectra of moment operators via mean-field techniques [71]. Reference [69] then rigorously showed that brickwork random circuits of depth $T\geq Ct^{9.5}\,[2Kt+\log_{2}(1/\varepsilon)]$ [^10] We use the parameter $T$ to denote the number of layers in a brickwork layout. This parameter is similar to the parameter $n$ presented in Sec. 4.1. This change in notation is to avoid confusion as $n$ is almost exclusively used for the system-size in the context of quantum information theory. are $\varepsilon$-approximate unitary $t$-designs, for a constant $C>0$ (whose form the authors calculated) and for all $t\leq O(2^{2K/5})$. An approximate design is a probability distribution that relaxes the equality in Eq. (14). In particular, the notion of approximation we get is surprisingly strong and implies relative errors (also called multiplicative errors) to Haar random unitaries for the outcome probabilities of any quantum experiment that queries the $t$ copies of $U$ in parallel. We refer to Refs. [69, 72] for details. This relaxation of the design property is sufficient to imply a circuit lower bound of $\mathcal{C}_{\delta}(U)\geq\Omega(Kt)$ with probability $1-e^{-\Omega(Kt)}$. In particular, the probability of significant short-cuts is small, not only in the system-size, but also in the number of copies $t$. The exponent $10.5$ of $t$ in this bound was improved to $5+o(1)$ [73]. Ref. [72] reduced the $t$ dependence to an optimal, linear scaling. More precisely, random circuits are $\varepsilon$-approximate unitary $t$-designs in depth
@@
$\mathcal{C}_{\delta}(U)$ denotes the minimal number of $2$-local unitaries necessary to generate any $V$ that is $\delta$-close to $U$ in operator-norm: $||U-V||\leq\delta$. Notice that the probability of significant short-cuts is negligible even in $T$. In particular, for very deep circuits, this becomes increasingly crucial as the number of random bits that are required to draw a circuit of depth $T$ scales like $\sim KT$.
@@
Similar techniques can be used to study the growth of quantum circuit complexity in models of time-dependent dynamics. For example, a discretized model of Brownian motion converges to approximate designs as quickly as random circuits do [75]. Reference [76] provides evidence for a similar convergence by continuous Brownian dynamics.
@@
4.6 Barriers to proving that quantum complexity grows superpolynomially under long-time evolutions
Each of the methods discussed above exploits a counting argument. The argument concerns a placeholder quantity that roughly measures the ensemble’s randomness. Finding such a quantity dependent on the ensemble is far easier than finding a similarly growing quantity for individual circuits. In the exact-circuit-complexity proof, the placeholder is the “accessible dimension” $\dim\mathcal{U}_{R}$. In the approximate-circuit-complexity proof, any quantity that signifies the convergence to designs (such as frame potentials [70] or the expected overlaps $\mathbb{E}_{V_{1},\ldots,V_{d}}|\langle\psi|V_{d}\cdots V_{1}|0^{K}\rangle|^{2}$ for random $2$-qubit gates $V_{1},\ldots,V_{d}$) will do). This strategy dominates all our proofs of lower bounds on circuit complexity beyond linear depth. Indeed, this limitation might be fundamental as circuit complexity is known to be notoriously difficult to bound.
But this pessimism concerns individual states. What about time evolution under randomly drawn local Hamiltonians? This scenario is closer to the setting considered by Brown and Susskind [44], who conjectured that the quantum complexity grows linearly for an exponentially long time under a typical time-independent local Hamiltonian. Still, the total number of distinct local time independent? Hamiltonians is upper bounded by $2^{\mathrm{poly}(K)}$, so clearly the above counting argument fails to establish superpolynomial lower bounds on the approximate quantum circuit complexity.
The problem of proving lower bounds on the quantum complexity for the time-evolution of local Hamiltonians turns out to be related to the separation of computational-complexity classes: Aaronson and Susskind [42] point out that the circuit complexity of exponentially long-time evolution under local Hamiltonians is superpolynomial if and only if $\mathrm{PSPACE}\not\subset\mathrm{BQP}/\mathrm{poly}$. $\mathrm{PSPACE}$ is the class of all computations performable with polynomial memory. $\mathrm{BQP}/\mathrm{poly}$ is the class of all problems solvable with (nonuniform) families of polynomially sized quantum circuits. Separations of complexity classes are notoriously difficult to prove, and few such separations are known. Proving a separation of the kind $\mathrm{PSPACE}\not\subset\mathrm{BQP}/\mathrm{poly}$ appears to require a breakthrough in theoretical computer science. Therefore, even showing the existence of a local Hamiltonian for which the complexity does not saturate after a polynomial time requires major technical advances.

### v2

@@ 4.2 Complexity saturation as a late-time stage of quantum many-body equilibration
The second law of thermodynamics suggests that state complexity should grow under generic dynamics. Of course, the second law concerns entropy, not complexity, and originally described classical systems. Yet complexity seems like it ought to obey the spirit of the second law [55]. In fact, there is an analogy, with a late time saturation of complexity characterizing the final stage of quantum many-body equilibration.
@@
To review the second law, we recall a classical monatomic gas in a box. Suppose that the gas particles are bunched together in a corner. The particles will spread across the box: the entropy of the gas, calculated with the Sackur-Tetrode equation, will end up greater than it was originally. The gas will equilibrate internally. Furthermore, suppose that the particles interact. They will thermalize; a particle’s phase-space distribution will approach the Maxwell–Boltzmann distribution.
@@
Given the second law’s ubiquity, one should expect generic quantum many-body systems to also equilibrate and thermalize internally. Quantum many-body systems undergo more stages of equilibration and thermalization than classical systems do. We will describe some stages, which end with complexity saturation. First, we discuss the kinds of systems that are expected to thermalize.
@@
Which systems thermalize is difficult to predict and prove. Thermalization is often attributed to chaos, but there are many definitions of quantum chaos [12, 13, 14]. According to early thinking, a system is quantum-chaotic if it has a chaotic semiclassical limit. Later studies defined chaos through energy-level statistics modeled by Wigner–Dyson distributions. Recent work has defined chaos through properties of out-of-time-ordered correlators [56] and spectral form factors [57, 58, 59, 12], reviewed below. Some authors acknowledge challenges of characterizing quantum chaos by writing “chaos” within quotation marks and taking the term to mean nonintegrability [60]. Yet debate surrounds the definition of quantum nonintegrability, too. According to a common definition, a system is integrable if it has extensively many nontrivial conserved quantities [13].777 Consider eigendecomposing an arbitrary quantum Hamiltonian: $H=\sum_{j}E_{j}\lvert j\rangle\!\langle j\rvert$. The Hamiltonian conserves every eigenprojector $\lvert j\rangle\!\langle j\rvert$. Furthermore, the number of eigenprojectors is exponentially large in the system’s size. Yet these facts do not render $H$ integrable; if they did, every $H$ would be integrable. A conserved quantity must at least not be a simple sum of eigenprojectors in order to be considered nontrivial. To compound the ambiguity, many equilibration results follow from empirical observations, rather than from proofs, and govern specific systems, rather than general ones. We will therefore write “chaotic” within quotation marks in this subsection to communicate expectations about behaviors common to many systems called chaotic or nonintegrable.
@@
In reviewing stages of quantum many-body equilibration, we refer to the following setup. Consider a closed quantum system of $K\gg 1$ subsystems—for example, qubits. Let the Hamiltonian $H$ be “chaotic”; and let the interactions be $k$-body, for some $k\in\{1,2,\ldots,K\}$. Suppose that the system begins in a simple, pure nonequilibrium state, such as an $K$-qubit product state.
@@
We now define two quantities—the out-of-time-ordered correlator and spectral form factor—used to define stages of equilibration. Let $W$ and $V$ denote local, Hermitian or unitary operators localized far apart. Examples include Pauli operators that act on distant qubits. Define the Heisenberg-picture operator $W(t)\coloneqq e^{iHt}We^{-iHt}$, and consider an arbitrary state $\rho$. The (four-point) out-of-time-ordered correlator is ${\rm Tr}\bm{(}W^{\dagger}(t)V^{\dagger}W(t)V\rho\bm{)}$; we discuss its significance below. Define the partition function $Z(\beta)\coloneqq{\rm Tr}(e^{-\beta H})$. Taking $\beta,t\in\mathbb{R}$, then the spectral form factor is $|Z(\beta+it)/Z(\beta)|^{2}$. It measures correlations across a Hamiltonian’s spectrum.
@@
    Local equilibration: Two-point correlators decay to $\sim 1/e$ of their initial values. Time-ordered correlators relax to approximately their long-time values. Small subsystems thermalize. This equilibration occurs at the dissipation time $t_{\rm D}$ [10], also called the thermalization time. The same timescale is called the collision time if a system admits a quasiparticle description. Under certain conditions (e.g., if the interactions are highly nonlocal), $t_{\rm D}$ is constant in $K$, the number of degrees of freedom. Under other conditions, $t_{\rm D}\sim K$. For example, this latter scaling characterizes a one-dimensional system with nearest-neighbor interactions, by the Lieb–Robinson bound [61].
@@
    Quantum information scrambling: Initially localized information spreads across the system through many-body entanglement. Out-of-time-ordered correlators decay from their $O(1)$ initial values [56]. This scrambling occurs around the scrambling time, $t_{*}$. $t_{*}$ can be parametrically larger than $t_{\rm D}$, for example if there are all-to-all couplings [52]. A lower bound limits how quickly many-body entanglement can form: $t_{*}\geq\frac{\hbar}{2\pi k_{\rm B}T}\,\log(K)$ [10]. This bound saturates only if $t_{\rm D}$ is constant. In the context of semiclassical, single-particle quantum chaos, an analogue of $t_{*}$ is called the Ehrenfest time [62].
@@
    Breakdown of quantum–classical correspondence and rise in predictiveness of random-matrix theory: This stage can begin around the same time as scrambling. Hence we have already seen one name for the time when this stage takes place: the Ehrenfest time $t_{\rm E}$, also called the Thouless time $t_{\rm Th}$. This time scale varies logarithmically with the system size: $t_{\rm E}\sim\log(K)$. How this stage manifests depends on the system and measures studied. An example involves transport in, e.g., electronic systems [63, 64]. Here, an initially narrow wave packet expands to cover a classically relevant length scale around $t_{\rm E}$. The quantum system’s behavior diverges from its classical analogue; hence the name Ehrenfest time alludes to Ehrenfest’s theorem, which highlights a parallel between quantum and classical systems. In another example, at this stage, the spectral form factor stops declining and begins to ramp upward again: random-matrix theory begins to accurately predict the spectral form factor, which ceases to depend on microscopic details [65, 66].
@@
    The spectral form factor changes from increasing linearly with time to remaining constant: This change happens at the Heisenberg time $t_{\rm H}$, or plateau time [65]. Denote by $\Delta$ the average inverse gap between consecutive eigenenergies. The Heisenberg time is $t_{\rm H}\coloneqq h/\Delta\sim e^{K}$ [64].
@@
    Complexity saturation: The state’s complexity grows to $\sim e^{K}$. The saturation time has been conjectured to scale as $e^{K}$. This conjecture has been proven under certain conditions (Sec. 4.3).
As a caveat: different authors use the same terms differently when referring to stages of quantum many-body equilibration. Also, the above list is intended to be illustrative, not comprehensive; compiling all the stages [14, 67, 68, 66, 3, 69] would require another review. For example, subtleties in the spectral form factor define other stages; see Sec. 1.1 of [67], and Sec. 1.4.3 of [66], for clear synopses and for references. Operator dynamics in Krylov space and the spread of wavefunctions require their own treatment [3, 21, 22]. Additionally, if $K$ is finite, then revivals (Poincaré recurrences and complexity revivals) occur long after times $\sim e^{K}$. Such revivals temporarily undo the equilibration but are consistent with the second law for systems of finite size.
@@ 4.3 Growth of complexity under random circuits
Random circuits are a toy model for chaotic quantum dynamics. We can prove random-circuit versions of conjectures originally formulated about the complexity of chaotic dynamics. When forming a random circuit, we choose the gates independently from a probability distribution on the unitary group, for example $\mathrm{SU}(4)$ if we have a four dimensional Hilbert space. Typically, we choose gates according to the Haar measure $\mu_{\rm H}$ on the unitary group, the unique distribution that is invariant under left- and right- multiplication by any unitary. $\mu_{\rm H}$ generalizes the uniform distribution on an interval, as well as the Lebesgue measure, to (locally compact) Lie groups. $\mu_{\rm H}$ can be viewed as the uniform measure on these groups.
Known barriers such as the natural proof barrier discovered by Razborov and Rudich [70, 71, 72] prevent us from proving superpolynomial lower bounds on the circuit complexity of an explicit Boolean function. The minimal number of elementary gates (such as AND, NAND, and XOR) in a Boolean function’s implementation is called its circuit complexity. A similar obstructions might inhibit us from proving superpolynomial lower bounds on the quantum complexity of quantum states, that is, the minimal number of $2$-local quantum gates required for construction of the target state from a similar reference. One can more easily prove lower bounds on the complexities of states prepared with random circuits as the latter quickly assume properties of the Haar measure, counting arguments about which yield lower bounds on the circuit complexity. In 1949 Claude Shannon [73] observed that the number of Boolean functions $f:\{0,1\}^{n}\to\{0,1\}$ outgrows the number of functions implementable with less than $\sim 2^{n}/n$ gates. A similar counting argument applies to Haar-random unitaries [74].
Sections 4.4 and Section 4.5 review two results about the growth of quantum circuit complexity. Both arguments followed from adapting Haar-measure counting arguments to random circuits. In a sense, these arguments offer to replace statements about vaguely defined “generic” dynamics reviewed in Section 4.1 with well-defined statements about average dynamics. These statements should apply to the dynamics of “most” systems and so should be “generic” in a well defined sense. We will arrive at these results by thinking about the statistics/distribution of unitaries and circuits, where “genericity” acquires the sense of being “very likely” or “with likelihood=1”. Section 4.6 reviews why we still lack satisfying lower bounds on the quantum complexity of chaotic dynamics.
@@ 4.4 Linear growth of exact circuit complexity
Here, we review a proof of the linear growth of the exact circuit complexity [75] for random circuits. More precisely, we show that the exact circuit complexity grows linearly with the number of layers, up to exponentially depth, with unit probability over the choice of circuits. Moreover, the complexity saturates for exponentially deep circuits, again with unit probability. The exact circuit complexity quantifies the minimal number of gates required to prepare a state or effect a unitary exactly. The preparation and action of the unitary cannot include errors. Consequently, some unitaries arbitrarily close to the identity operator have exact circuit complexities of $e^{\Omega(K)}$, [^8] In this section, we use big-Omega notation, $\Omega(X)$, to mean, roughly, “a term that grows at least as quickly as $X$”. where $K$ denotes the system size, or number of degrees of freedom (qubits).
We now sketch a proof of the conjecture [55] that circuit complexity grows linearly for an exponentially long time, as applied to random circuits and the exact circuit complexity. We will compare the degrees of freedom in two sets of unitaries: the unitaries generated by the depth-$D$ circuits and the unitaries generated by smaller, depth-$D^{\prime}$ circuits. The basic idea echoes the fact that, in $\mathbb{R}^{3}$, a plane has no volume. In terms of a well-behaved probability measure on a manifold $M$, every lower-dimensional submanifold has a probability of 00. Denote by $\mathcal{U}_{R}$ the set of all unitaries that can be generated by $R$ gates each. Fortunately, $\mathcal{U}_{R}$, although not a manifold, turns out to be a semialgebraic set defined by constraints $\{f_{i}\geq 0,g_{i}>0,h_{i}=0\}$, where $f_{i},g_{i},h_{i}:{\rm SU}(D)\to\mathbb{R}$ are polynomials. Every semialgebraic set has a well-defined dimension similar to a manifold’s dimensionality.
@@ 4.5 Linear growth of approximate circuit complexity
We can combine this bound with the quick convergence of random circuits to unitary designs. Brown and Viola provided evidence for fast convergence by analyzing the spectra of moment operators via mean-field techniques [83]. Reference [81] then rigorously showed that brickwork random circuits of depth $T\geq Ct^{9.5}\,[2Kt+\log_{2}(1/\varepsilon)]$ [^9] We use the parameter $T$ to denote the number of layers in a brickwork layout. This parameter is similar to the parameter $n$ presented in Sec. 4.1. This change in notation is to avoid confusion as $n$ is almost exclusively used for the system size in the context of quantum information theory. are $\varepsilon$-approximate unitary $t$-designs, for a constant $C>0$ (whose form the authors calculated) and for all $t\leq O(2^{2K/5})$. An approximate design is a probability distribution that relaxes the equality in Eq. (14). In particular, the notion of approximation we get is surprisingly strong and implies relative errors (also called multiplicative errors) to Haar random unitaries for the outcome probabilities of any quantum experiment that queries the $t$ copies of $U$ in parallel. See Refs. [81, 84] for details. This relaxation of the design property is sufficient to imply a circuit lower bound of $\mathcal{C}_{\delta}(U)\geq\Omega(Kt)$ with probability $1-e^{-\Omega(Kt)}$. In particular, the probability of significant short-cuts is small, not only in the system-size, but also in the number of copies $t$. The exponent $10.5$ of $t$ in this bound was improved to $5+o(1)$ in [85] and Ref. [84] reduced the $t$ dependence to an optimal, linear scaling. More precisely, random circuits are $\varepsilon$-approximate unitary $t$-designs in depth
@@
$\mathcal{C}_{\delta}(U)$ denotes the minimal number of $2$-local unitaries necessary to generate any $V$ that is $\delta$-close to $U$ in operator-norm: $||U-V||\leq\delta$. Notice that the probability of significant short-cuts is negligible. In particular, for very deep circuits, this becomes increasingly crucial as the number of random bits that are required to draw a circuit of depth $T$ scales like $\sim KT$.
@@
Similar techniques can be used to study the growth of quantum circuit complexity in models of time-dependent dynamics. For example, a discretized model of Brownian motion converges to approximate designs as quickly as random circuits do [87]. Reference [88, 89] provides evidence for a similar convergence by continuous Brownian dynamics. These continuous versions can be directly formulated in the context of holography [90, 91] as we discuss in Sec. 7.
@@
4.6 Barriers to proving that quantum complexity grows superpolynomially at late times
Each of the methods discussed above exploits a counting argument. The argument concerns a placeholder quantity that roughly measures the ensemble’s randomness. Finding such a quantity for an ensemble is far easier than finding a similarly growing quantity for individual circuits. In the exact-circuit-complexity proof, the placeholder is the “accessible dimension” $\dim\mathcal{U}_{R}$. In the approximate-circuit-complexity proof, any quantity that signifies the convergence to designs will do, such as frame potentials [82] or the expected overlaps $\mathbb{E}_{V_{1},\ldots,V_{d}}|\langle\psi|V_{d}\cdots V_{1}|0^{K}\rangle|^{2}$ for random $2$-qubit gates $V_{1},\ldots,V_{d}$. This strategy dominates all our proofs of lower bounds on circuit complexity beyond linear depth. Indeed, this limitation might be fundamental as circuit complexity is known to be notoriously difficult to bound.
But these comments apply to individual states. What about time evolution under randomly drawn local Hamiltonians? This scenario is closer to the setting considered by Brown and Susskind [55], who conjectured that the quantum complexity grows linearly for an exponentially long time for a typical time-independent local Hamiltonian. Still, the total number of distinct local time independent Hamiltonians is upper bounded by $2^{\mathrm{poly}(K)}$; so clearly the above counting argument fails to establish superpolynomial lower bounds on the approximate quantum circuit complexity.
The problem of proving lower bounds on the quantum complexity for the time-evolution of local Hamiltonians turns out to be related to the separation of computational-complexity classes: Aaronson and Susskind [53] point out that the circuit complexity of exponentially long-time evolution by local Hamiltonians is superpolynomial if and only if $\mathrm{PSPACE}\not\subset\mathrm{BQP}/\mathrm{poly}$. $\mathrm{PSPACE}$ is the class of all computations that can be performed with polynomial memory. $\mathrm{BQP}/\mathrm{poly}$ is the class of all problems that can be solved with (nonuniform) families of polynomially sized quantum circuits. Separations of complexity classes are notoriously difficult to prove, and few such separations are known. Proving a separation of the kind $\mathrm{PSPACE}\not\subset\mathrm{BQP}/\mathrm{poly}$ appears to require a breakthrough in theoretical computer science. Therefore, even showing the existence of a local Hamiltonian for which the circuit complexity does not saturate after a polynomial time requires major technical advances.